OpenStudy (anonymous):

a group of 625 students has a mean age of 15.8 years with a standard deviation of 0.6 years. the ages re normally distributed. how many students are younger than 16.2 years?

OpenStudy (anonymous):

Normal distribution means we can us z-scores \(z=\frac{x-\mu}{\sigma}=\frac{16.2-15.8}{0.6}=0.667\) Find the area to the left of 0.667 in the z-table for your answer.